# GeV-scale QCD Axion

### v1

@@ XII Astrophysical Constraints
The only astrophysical constraint for a heavy axion is from SN1987a, which applies to a mass up to 500 MeV [lee2018revisitingsupernova1987alimits] if the dominant coupling is to photons. However, our axion couples to nucleons with the Yukawa coupling Eq. (41) and the axion is trapped, not leading to cooling of the proto-neutron star core. We conclude that there are no astrophysical constraints.
@@ XIII A UV Completion
An ultraviolet (UV) completion of the model has two scalar multiplets beyond the standard model: an additional Higgs doublet $\Phi$ and a complex scalar $\phi$ neutral under the standard-model gauge groups. Both of them carry $U(1)_{PQ}$ charge +1+1, while $u_{R}$ 1-1.
The most general Yukawa couplings to quarks are
${\cal L}_{Y}$ $=-Y_{u}^{ij}\bar{Q}_{i}U_{j}H-Y_{d}^{ij}\bar{Q}_{i}D_{j}\tilde{H}-\kappa^{i}\bar{Q}_{i}u_{R}\Phi+h.c.$ (43)
where $\tilde{H}=i\sigma_{2}H^{*}$. $U_{j}=(t_{R},c_{R})^{T}$ have only two components because of the Peccei–Quinn charge. Using the usual bi-unitary transformations, we can make $Y_{u}^{ij}$ real, positive, and diagonal, and $Y_{d}^{ij}=V_{ij}Y_{d_{i}}^{\rm diag}$ where $Y_{d_{i}}^{\rm diag}$ is real, positive, and diagonal, $V_{ij}$ the usual Cabibbo–Kobayashi–Maskawa (CKM) matrix, and $\kappa^{i}$ are all real and positive. It can be written as
${\cal L}_{Y}=$ $-(\bar{Q}_{t},\bar{Q}_{c})\left(\begin{array}[]{c}y_{t}t_{R}\\ y_{c}c_{R}\end{array}\right)H-(\bar{Q}_{t},\bar{Q}_{c},\bar{Q}_{u})V\left(\begin{array}[]{c}y_{b}b_{R}\\ y_{s}s_{R}\\ y_{d}d_{R}\end{array}\right)\tilde{H}$ (49)
$-(\kappa_{t}\bar{Q}_{t}+\kappa_{c}\bar{Q}_{c}+\kappa_{u}\bar{Q}_{u})u_{R}\Phi+h.c.$ (50)
without a loss of generality. Here, $\tilde{H}=i\sigma_{2}H^{*}$, $y_{q}$ real and positive, and $V_{ij}$ the CKM matrix, and $\kappa_{i}$ are all real and positive.
The most general scalar potential is
$V=$ $\lambda_{1}(H^{\dagger}H-v^{2})^{2}+M_{\Phi}^{2}\Phi^{\dagger}\Phi+m_{\phi}^{2}\phi^{*}\phi+\lambda_{2}(\Phi^{\dagger}\Phi)^{2}$
$+\lambda_{3}(\phi^{*}\phi)^{2}+\lambda_{4}H^{\dagger}H\Phi^{\dagger}\Phi+\lambda_{5}H^{\dagger}H\phi^{*}\phi$
$+\lambda_{6}\Phi^{\dagger}\Phi\phi^{*}\phi+\mu_{\Phi}(\Phi^{\dagger}H\phi+c.c.).$ (51)
We assume $M_{\Phi}^{2}>0$ and no expectation value of $\Phi$ at the electroweak scale. Once $H$ acquires an expectation value, the neutral component $\Phi^{0}$ and $\phi$ mix,
$V\supset(\Phi^{0*},\phi^{*})\left(\begin{array}[]{cc}M_{\Phi}^{2}&\mu_{\Phi}v\\ \mu_{\Phi}v&m_{\phi}^{2}\end{array}\right)\left(\begin{array}[]{c}\Phi^{0}\\ \phi\end{array}\right).$ (56)
We assume that $M_{\Phi}$ is above 5 TeV as discussed below, while the lighter eigenstate $\varphi=\phi\cos\vartheta-\Phi^{0}\sin\vartheta$ has a mass $m_{\varphi}$ in the GeV range. The mixing angle is $\vartheta\approx\arctan(\mu_{\Phi}v/M_{\Phi}^{2})$. This requires a fine-tuning of order $({\rm GeV}/5~{\rm TeV})^{2}\sim 10^{-7}$. Note that the DFSZ axion models require a much more precise fine-tuning of order $(v/f_{a})^{2}<10^{-12}$ for $f_{a}>5\times 10^{8}$ GeV given the pulsar constraints [Hamaguchi:2018oqw]. This fine-tuning is clearly related to the hierarchy problem which may be ameliorated by supersymmetry. Further model building efforts along this line are beyond the scope of this paper.
After integrating out massive fields, the Lagrangian has additional couplings
${\cal L}_{QCD}\supset-(\kappa^{\prime}_{t}\varphi\,\overline{t_{L}}u_{R}+\kappa^{\prime}_{c}\varphi\,\overline{c_{L}}u_{R}+c.c.).$ (57)
Here, $\kappa^{\prime}_{i}\approx\kappa_{i}\sin\vartheta$ are the couplings induced by the mixing between $\Phi^{0}$ and $\varphi$, and $\kappa=\kappa_{u}\sin\vartheta$ in Eq. (4).
XIV More Constraints
The UV completion discussed above has additional experimental constraints.
We see that there is generically a coupling $\kappa^{\prime}_{c}\varphi\,\overline{c_{L}}u_{R}$. Without a $\lambda^{\prime}(\varphi^{*}\varphi)^{2}$ term in the potential where $\lambda^{\prime}=\lambda_{3}\sin^{4}\vartheta$, it does not lead to the $D^{0}$$\overline{D}^{0}$ mixing. Once the term is included, the degeneracy between $\sigma$ and $a$ is broken, and it can induce the $D^{0}$$\overline{D}^{0}$ mixing. We simply have to assume that
$\lambda^{\prime}\kappa^{\prime}_{c}<\lambda^{\prime}\left(m^{2}_{\varphi}\frac{1}{4\pi^{2}}G_{F}^{2}\left|V_{cs}V_{cd}\right|^{2}m_{s}^{2}\right)^{1/2}\approx 5\times 10^{-8}.$ (58)
We do not need this coupling for our model and simply assume it is small enough. Note that the structure of Yukawa couplings is one of the mysteries in the standard model and requires a better understanding of the origin of flavor. The coupling $\lambda^{\prime}$ is not needed either and it could well be one-loop suppressed.
Given that $\varphi$ has an admixture of electroweak doublet $\Phi$, it can be produced in $Z$ decay, $Z\rightarrow\sigma a$. The branching fraction is approximately
${\rm BR}(Z\rightarrow\sigma a)=\frac{1}{6}{\rm BR}(Z\rightarrow\nu\bar{\nu})\sin^{4}\vartheta=0.033\sin^{4}\vartheta.$ (59)
Since both $\sigma$ and $a$ dominantly decay into multi-pion states with a high Lorentz boost, it would appear as a two-jet event. The hadronic width has an excellent agreement between theory and experiment within the error at the level of $\Delta{\rm BR}(Z\rightarrow{\rm hadrons})=0.056\%$. We therefore need $\vartheta<0.37$. Assuming $\kappa_{u}\lesssim 1$ to be perturbative, we find $\kappa=\kappa_{u}\sin\vartheta\lesssim 0.37$. Note that $\kappa$ goes up from $m_{Z}$ to 2 GeV from QCD renormalization group equations and the limit is approximately $\kappa\lesssim 0.5$ as mentioned earlier. It is interesting that the extracted $\alpha_{s}(m_{Z})$ from the $\Gamma(Z\rightarrow{\rm hadrons})$ is $\alpha_{s}(m_{Z})=0.1224\pm 0.0038$ [ParticleDataGroup:2004fcd], larger than the current PDG average of $\alpha_{s}(m_{Z})=0.1180(9)$, which may imply additional hadronic final states in the $Z$ decay.
Another constraint is the LHC search for a dijet resonance. In our model, there are resonant $\bar{d}\rightarrow\Phi^{+}\rightarrow u\bar{d}, uu$$\bar{u}\rightarrow\Phi^{0}\rightarrow u\bar{u}$, and $d\bar{u}\rightarrow\Phi^{-}\rightarrow d\bar{u}$ productions. The experimental limits [ATLAS:2019fgd, CMS:2019gwf] can be re-interpreted as a lower bound $M_{\Phi}\gtrsim 5$ TeV.
XV Conclusion
@@
It is rather surprising that the Peccei–Quinn breaking scale may be below the QCD scale. In this case, the axion and its scalar partner are at the GeV range, and it is tantalizing to think that they may be hiding among the observed hadronic resonances. Given the high mass, it is immune to quantum gravity corrections. In a UV completed model, dijet resonance at muli-TeV can be searched for at the LHC. The mass splitting between $\pi^{\pm}$ and $\pi^{0}$ is the most significant constraint.
XVI Acknowledgments
@@ Acknowledgements.
HM thanks Tsutomu Yanagida and Ben Safdi for useful comments. This work was supported by the NSF grant PHY-2515115, by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under the Contract No. DE-AC02-05CH11231, by the JSPS Grant-in-Aid for Scientific Research JP23K03382, Hamamatsu Photonics, K.K, Tokyo Dome Corporation, and by the World Premier International Research Center Initiative (WPI) MEXT, Japan.

### v2

@@ XII Astrophysical Constraints
The only astrophysical constraint for a heavy axion is from SN1987a, which applies to a mass up to 500 MeV [35] if the dominant coupling is to photons. However, our axion couples to nucleons with the Yukawa coupling Eq. (43) and the axion is trapped, not leading to cooling of the proto-neutron star core. We conclude that there are no astrophysical constraints.
@@ XIII A UV Completion
We need a dimension-five operator
$\frac{y\kappa_{u}}{M_{U}}\bar{Q}_{u}\varphi Hu_{R}+c.c.$ (45)
for the coupling $\kappa$ in Eq. (4), with $\langle H\rangle=v=174$ GeV,
$x\kappa$ $=\frac{yv}{M_{\cal Q}}\kappa_{u}.$ (46)
A UV completion of the operator Eq. (45) has a heavy vector-like quark ${\cal Q}=({\cal U},{\cal D})$ with the same quantum numbers as the left-handed quarks except for a $U(1)_{PQ}$ charge 1-1. The most general couplings are
${\cal L}_{U}=$ $\bar{\cal Q}(i{\not\!\!D}-M_{\cal Q}){\cal Q}+(\kappa_{i}\bar{Q}_{i}\varphi{\cal Q}+y\bar{\cal Q}Hu_{R}+c.c.)$ (47)
Integrating out $\cal Q$}, we indeed obtain Eq. (45). The vector-like quark can be produced in pairs at the LHC. The best limit $M_{\cal D}>1.53$ TeV comes from the pair production of $\cal D$} followed by ${\cal D}\rightarrow u_{R}W^{-}$ or its CP conjugate [1]. With $y,\kappa_{u}\approx 1$ and $M_{\cal Q}\approx 1.6$ TeV, we find $\kappa\approx 0.11$, which is enhanced to $\kappa\approx 0.25$ down to the QCD scale, sufficiently large for our purposes. There is also a limit on $\cal U$} that decays into $u_{L}+\varphi$, which appear as four-jet events with di-jet resonances, limiting $M_{\cal U}>770$ GeV [50] even though a few holes exist. A dedicated search for ${\cal U}\rightarrow uh$ would be very interesting. Single production searches at the LHC assume the $tW$ or $bW$ final states [49] and do not apply. For the light quark final states, the limits go back to the Tevatron experiments [2] that limit masses up to 700 GeV at most.
The light up-quark eigenstates have a mixing between ${\cal U}_{L}$ and $u_{L}$ only by $\kappa_{u}\langle\varphi\rangle/M_{\cal Q}\simeq 10^{-3}$. Therefore there is practically no modifications to the charged-current weak interactions and the impact on the CKM unitarity is only at the level of $10^{-6}$. On the other hand, the neutral-current weak interactions are affected by the large mixing between ${\cal U}_{R}$ and $u_{R}$, and hence the $Z$ coupling to the light right-handed up-quark eigenstate is modified as
$-\frac{2}{3}\sin^{2}\theta_{W}\rightarrow-\frac{2}{3}\sin^{2}\theta_{W}+\frac{1}{2}\frac{(yv)^{2}}{M_{\cal Q}^{2}+(yv)^{2}}\ .$ (48)
For $M_{\cal Q}=1.6$ TeV and $y=1$, it modifies ${\rm BR}(Z\rightarrow u_{R}\overline{u_{R}})$ by $0.15\%$, well within the measurement error of $\Delta{\rm BR}(Z\rightarrow u\bar{u}+c\bar{c})=0.6~\%$. On the other hand, the impact on the hadronic width is nominally at the $3\sigma$ level compared to the measurement error, $\Delta{\rm BR}(Z\rightarrow{\rm hadrons})=0.06~\%$. It will surely be tested by future Giga-$Z$ or Tera-$Z$ experiments. It is also subject to a question of contribution to the $Z$ decay into down-type quarks from new physics, beyond the scope of this work.
The dimension-five operator (45) leads to a decay
$\Gamma(h\rightarrow\overline{u_{L}}u_{R}\varphi,\overline{u_{R}}u_{L}\varphi^{*})$
$=\left(\frac{y\kappa_{u}}{M_{U}}\right)^{2}\frac{m_{h}^{3}}{512\pi^{3}}=0.21~{\rm MeV}\left(\frac{y\kappa_{u}770~{\rm GeV}}{M_{U}}\right)^{2}.$ (49)
independent of the UV completion. It is comparable with the predicted $\Gamma(h\rightarrow gg)=0.335$ MeV [40]. At the LHC, this decay mode would not be observable. At a future $e^{+}e^{-}$ Higgs factory, the measurement of this decay mode can be probed as a part of the $\Gamma(h\rightarrow{\rm hadrons})$ measurement with the bottom and charm veto.
The coupling $\kappa_{c}$ can lead to $D^{0}$$\overline{D}^{0}$ mixing at the one-loop level with a box diagram of $\cal U$} and $\varphi$. This coupling is not needed in our proposal and we simply assume it is small.
XIV Conclusion
@@
It is rather surprising that the Peccei–Quinn breaking scale may be below the QCD scale. In this case, the axion and its scalar partner are at the GeV range, and it is tantalizing to think that they may be hiding among the observed hadronic resonances. Given the high mass, it is immune to quantum gravity corrections. In a UV completed model, dijet resonance in four-jet events can be searched for at the LHC, and an enhanced $\Gamma(h\rightarrow gg)$ at a future $e^{+}e^{-}$ Higgs factory. The mass splitting between $\pi^{\pm}$ and $\pi^{0}$ is the most significant constraint.
XV Acknowledgments
@@ Acknowledgements.
HM thanks Tsutomu Yanagida, Ben Safdi, and Gilad Perez for useful comments. Special thanks go to Satoshi Shirai who checked aspects of UV completion. This work was supported by the NSF grant PHY-2515115, by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under the Contract No. DE-AC02-05CH11231, by the JSPS Grant-in-Aid for Scientific Research JP23K03382, Hamamatsu Photonics, K.K, Tokyo Dome Corporation, and by the World Premier International Research Center Initiative (WPI) MEXT, Japan.